How many grams of oxygen gas would have been required to produce 1000 grams of nitric acid HNO3?
1
Expert's answer
2021-09-29T05:03:04-0400

4NH3 + 5O2 → 4NO + 6H2O (1)

2NO + O2 → 2NO2 (2)

3NO2 + H2O → 2HNO3 + NO (3)

M(HNO3) = 63.01 g/mol

n(HNO3) "= \frac{m}{M} = \frac{1000}{63.01} = 15.87 \;mol"

According to the reaction (3):

n(NO2) = "\frac{3}{2}" n(HNO3) "= \frac{3}{2} \times 15.87 = 23.80 \;mol"

According to the reaction (2):

n(NO) = n(NO2) = 23.80 mol

According to the reaction (1):

n(O2) = "\frac{5}{4}" n(NO) "= \frac{5}{4} \times 23.80 = 29.75 \;mol"

M(O2) = 15.99 g/mol

m(O2) "= 15.99 \times 29.75 = 475.70 \;g"

Answer: 475.70 g